in an experiment, the following readings where observed; volume of alcohol flow pair minutes is 10^-5m^3, the density of alcohol is 800kg\m^3, length of tube is 0.5m, radius of tube is 0.05cm and the height of alcohol is 60cm. calculate the coefficient of viscosity of alcohol?
1
Expert's answer
2021-07-12T12:11:08-0400

"V=10^{-5}m^3min^{-1}=\frac{10^{-5}}{60}m^3s^{-1}\\
\rho=800kgm^{-3}\\
l=0.5m\\
r=0.05cm=5\times{10^{-4}}m\\
h=60cm=0.6m\\"

Pressure P is

"P=h\rho{g}\\
g=10ms^{-2}\\
P=10\times{0.6}\times{800}=4800Nm^{-2}\\"

Coefficient of viscosity is given by Poiseuille's formula

"\eta=\dfrac{\pi{P}r^4}{8Vl}\\
\eta=\dfrac{\pi{4800}(5\times10^{-4})^4}{8\times\frac{10^{-5}}{60}\times0.5}\\
\eta=1.41\times10^{-3}Nsm^{-2}"